# A Temporal-Spectral Fusion Transformer with Subject-specific Adapter for Enhancing RSVP-BCI Decoding

### v1

@@ 3.3 Data Acquisition and Preprocessing
Figure 4: The structure of our proposed TSformer-SA. The inputs consist of EEG temporal signals representing the temporal view and the spectrogram images representing the spectral view. The feature extractor tokenizes the inputs and extracts the view-specific features. Subsequently, the cross-view interaction module extracts the common features from both views and the fusion module fuses the two-view features for classification. The above three modules are trained during the pre-training stage and only the subject-specific adapter is trained in the fine-tuning stage.
@@ 4 Method
In this section, we introduce the Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) for RSVP decoding. As shown in Fig. 4, TSformer-SA is a symmetrical dual-stream Transformer comprising a feature extractor, a cross-view interaction module, a fusion module, and a subject-specific adapter. The inputs of the model are EEG temporal signals as the temporal view and the spectrogram of each channel computed using CWT as the spectral view. First, the feature extractor employs a slice embedding layer to tokenize the two-view input and an encoder layer to extract view-specific features. Subsequently, the cross-view interaction module employs cross-attention and token fusion mechanisms to realize information transfer between the two views. Simultaneously, we introduce a multi-view consistency loss for the common feature extraction between two views by maximizing the similarity of different view features from the same sample. Then, an attention-based fusion module achieves the feature fusion between two views to obtain comprehensive features for classification. The optimization strategy of TSformer-SA consists of the pre-training stage and the adapter-based fine-tuning stage. In the pre-training stage, multi-view consistency loss and cross-entropy loss are used to optimize TSformer on data from existing subjects. Then the proposed subject-specific adapter is fine-tuned on new subjects’ data to rapidly transfer the learned general patterns of the pre-trained model to the data distribution of new subjects.
@@ 4.1 Feature Extractor
To represent EEG temporal signals and spectrogram images consistently and realize global feature extraction within each view, we introduce a feature extractor based on the patch embedding layer and encoder layer in the Vision Transformer (ViT) dosovitskiy2020image . The feature extractor includes a slice embedding layer and an encoder layer with shared parameters across two streams.
@@
The inputs of the slice embedding layer are EEG temporal signals ($\boldsymbol{S}_{tem}\in\mathbb{R}^{1\times C\times T}$) and spectrogram images of each channel ($\boldsymbol{S}_{spe}\in\mathbb{R}^{S\times C\times T}$), where $C$ denotes the number of channels, $T$ denotes the number of sampling points, and $S$ denotes the size of scale in CWT. To tokenize EEG temporal signals, we slice and reshape the input signals $\boldsymbol{S}_{tem}$ into a sequence of flattened slices $\left[\boldsymbol{s}_{1}^{tem},\boldsymbol{s}_{2}^{tem},\cdots,\boldsymbol{s}_{n}^{tem}\right]$ along the time dimension. Each $\boldsymbol{s}_{i}^{tem}\in\mathbb{R}^{(C\times t)}$ represents brain activities over a period of time, where $t$ is the slice length and $n=[T/t]$ is the number of slices. Then, the slice embedding layer maps them to $d$ dimensions, where the output of these temporal slice embeddings are referred to as temporal tokens $\boldsymbol{X}_{tem}\in\mathbb{R}^{n\times d}$. The process can be formulated as follows:
$\boldsymbol{X}_{tem}=\left[\boldsymbol{s}_{1}^{tem},\boldsymbol{s}_{2}^{tem},\cdots,\boldsymbol{s}_{n}^{tem}\right]^{T}\boldsymbol{W}_{pos}+\boldsymbol{E}_{pos},$ (1)
where $\boldsymbol{W}_{pos}\in\mathbb{R}^{(C\times t)\times d}$ is the parameter matrix and $\boldsymbol{E}_{pos}\in\mathbb{R}^{n\times d}$ is the learnable positional parameters added to the slice embeddings to retain positional information. The resulting sequence of embedding vectors serves as the input to the encoder layers. For the spectrogram images, we apply a similar slicing operation along the time dimension and reshape them into a sequence $\left[\boldsymbol{s}_{1}^{spe},\boldsymbol{s}_{2}^{spe},\cdots,\boldsymbol{s}_{n}^{spe}\right]$, where each $\boldsymbol{s}_{i}^{spe}\in\mathbb{R}^{(S\times C\times t)}$ represents brain spectral activities over a period of time. The subsequent steps are the same as those used for the EEG temporal signals and the output of these spectrogram slice embeddings are referred to as spectral tokens. The tokens at the same positions in token sequences of both views capture the respective temporal and spectral information of the EEG signals within the same time period.
@@
The encoder layer comprises two sub-layers: a Multi-Head Self-Attention (MHSA) layer capturing global temporal dependencies within input tokens, and a position-wise Feed-Forward Network (FFN) layer with a hidden layer consisting of $h\times d$ dimensions extracting feature representations, where $h$ denotes the number of heads. All sub-layers utilize a residual connection and layer normalization to enhance the scalability of the model. Additionally, we introduce a skip connection between the beginning and end of the encoder layer. Finally, to reduce model parameters and enable the subsequent view interaction more reasonable and feasible, we share the encoder parameters between the two streams to project the features into the same feature space.
@@ 4.2.1 Cross-view Interaction Module
The cross-view interaction module consists of a cross-attention layer and a token fusion layer. The cross-attention layer achieves mutually common feature extraction between two views based on the similarity between their tokens. The token fusion layer allows each view to extract complementary information from the other to enhance feature representation.
@@
The cross-attention layer replaces the self-attention mechanism in the encoder layer with a cross-attention mechanism, while the rest of the structure remains identical. The input of the cross-attention layer are token sequences of two views: $\boldsymbol{X}_{tem}=[\boldsymbol{x}^{tem}_{1},\boldsymbol{x}^{tem}_{2},\cdots,\boldsymbol{x}^{tem}_{n}]^{T}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{X}_{spe}=[\boldsymbol{x}^{spe}_{1},\boldsymbol{x}^{spe}_{2},\cdots,\boldsymbol{x}^{spe}_{n}]^{T}\in\mathbb{R}^{n\times d}$, where $\boldsymbol{x}^{tem}_{i}\in\mathbb{R}^{d}$ and $\boldsymbol{x}^{spe}_{i}\in\mathbb{R}^{d}$ denote the $i$-th token in temporal and spectral token sequences, respectively. In the cross-attention layer, both views mutually extract common features from each other. To illustrate the cross-attention mechanism, we take the process of $\boldsymbol{X}_{tem}$ extracting features from $\boldsymbol{X}_{spe}$ as an example. The cross-attention mechanism is composed of a linear projection layer and an attention layer. Firstly, the linear projection layer maps input token sequences from two different views to three distinct sequential vectors: query ($\boldsymbol{Q}$), key ($\boldsymbol{K}$), and value ($\boldsymbol{V}$). These vectors are generated as follows:
$\boldsymbol{Q}=\boldsymbol{X}_{tem}\boldsymbol{W}_{Q},\quad\boldsymbol{K}=\boldsymbol{X}_{spe}\boldsymbol{W}_{K},\quad\boldsymbol{V}=\boldsymbol{X}_{spe}\boldsymbol{W}_{V},$ (2)
where linear matrices $\boldsymbol{W}_{Q}\in\mathbb{R}^{d\times d_{k}}$, $\boldsymbol{W}_{K}\in\mathbb{R}^{d\times d_{k}}$, and $\boldsymbol{W}_{V}\in\mathbb{R}^{d\times d_{v}}$ are learnable parameter matrices. The $d_{k}$ denotes the dimensions of the query ($\boldsymbol{Q}$) and key ($\boldsymbol{K}$), while $d_{v}$ denotes the dimension of the value ($\boldsymbol{V}$). Then the cross-attention mechanism can be formulated as follows:
$Attention(\boldsymbol{X}_{tem},\boldsymbol{X}_{spe})$ $=\sigma(\frac{\boldsymbol{Q}\boldsymbol{K}^{T}}{\sqrt{d_{k}}})\boldsymbol{V}$ (3)
$=\sigma(\frac{\boldsymbol{X}_{tem}\boldsymbol{W}_{Q}\boldsymbol{W}_{K}^{T}\boldsymbol{X}_{spe}}{\sqrt{d_{k}}})\boldsymbol{X}_{spe}\boldsymbol{W}_{V},$
where $\sigma(\cdot)$ represents the softmax function. The mechanism is referred to as the self-attention mechanism when the query ($\boldsymbol{Q}$), key ($\boldsymbol{K}$), and value ($\boldsymbol{V}$) are all projected from the same token sequence. To overcome the constraint of limited modeling capacity in a single-head attention block, we employ a multi-head attention mechanism. This mechanism linearly projects the input into multiple feature subspaces with varying dimensions and processes them independently through parallel attention heads. The resulting vectors are concatenated and then mapped to the final output with a dimension of $d$.
To enhance the representation of each view and reduce feature divergence between two views, we introduce a token fusion wang2022multimodal operation within the cross-view interaction module. The token fusion dynamically detects uninformative tokens in each view and enhances them by aggregating informative features from the other view. As an example, for the updating of temporal tokens $\boldsymbol{X}_{tem}$, we assess the importance of token $\boldsymbol{x}^{tem}_{i}$ using a scoring function denoted as $f(\cdot)$:
@@
$f(\boldsymbol{x}^{tem}_{i})=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sigma\left(\dfrac{\boldsymbol{X}_{tem}\boldsymbol{W}_{score}\boldsymbol{X}_{tem}^{T}}{\sqrt{d}}\right)_{(j,i)},$ (4)
@@
where $\boldsymbol{W}_{score}\in\mathbb{R}^{d\times d}$ is a learnable parameter matrix. The scoring function represents the accumulation of attention weights from other tokens towards $\boldsymbol{x}^{tem}_{i}$ in the self-attention matrix of temporal tokens $\boldsymbol{X}_{tem}$. Tokens that receive higher attention can be considered to contain more information. On the basis of $f(\cdot)$, we can establish the scoring function of temporal tokens as $F(\cdot)$, which can be formulated as:
$F(\boldsymbol{X}_{tem})=\left[f(\boldsymbol{x}^{tem}_{1})\cdot\boldsymbol{1},f(\boldsymbol{x}^{tem}_{2})\cdot\boldsymbol{1},\cdots,f(\boldsymbol{x}^{tem}_{n})\cdot\boldsymbol{1}\right],$ (5)
@@
$\boldsymbol{X}_{tem}=\dfrac{1}{2}[\boldsymbol{X}_{tem}+\boldsymbol{X}_{spe}\odot\mathbb{I}$ $(F(\boldsymbol{X}_{tem})<\theta_{tem})$ (6)
$\odot\mathbb{I}(F(\boldsymbol{X}_{spe})>\theta_{spe})],$
where $\mathbb{I}(\cdot)$ is an indicator asserting the subscript condition, which outputs a mask tensor with element 0 or 1. The threshold $\theta_{tem}$ and $\theta_{spe}$ are equal to the median of $F(\boldsymbol{X}_{tem})$ and $F(\boldsymbol{X}_{spe})$, respectively. The operator $\odot$ denotes the element-wise multiplication.
Finally, we use a self-attention layer to extract the global features of tokens with added supplementary information. Token fusion applies an identical operation to both temporal and spectral tokens. Consequently, both views obtain supplementary information from the other view.
@@ 4.2.2 Multi-view Consistency Loss
In multi-view interactive learning, we introduce a multi-view consistency loss to maximize the similarity between two views features of the same sample extracted by the cross-view interaction module. Multi-view consistency loss is based on contrastive learning chen2020simple and contributes to facilitating effective common feature extraction between the two views. We use a convolution layer with four convolution kernels, each of size $(16,d/8)$ and stride $(16,d/8)$, to map the token sequences of both views to features $\{\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i}\}_{i=1}^{N}$ in a shared feature space, where $\boldsymbol{z}^{tem}_{i}$ and $\boldsymbol{z}^{spe}_{i}$ are the feature vectors of the temporal view and spectral view from the $i$-th sample, respectively. The $N$ denotes the batch size. Feature similarity is assessed using the cosine distance $sim(\boldsymbol{u},\boldsymbol{v})=\boldsymbol{u}^{T}\boldsymbol{v}/\|\boldsymbol{u}\|_{2}\|\boldsymbol{v}\|_{2}$, and the loss function with respect to positive pairs $(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})$ is defined as:
$\mathcal{L}$ =${}_{i}=\log\frac{e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})/\tau}}{\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{tem}_{j})/\tau}+\sum_{j=1}^{N}e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{j})/\tau}}$ (7)
$+\log\frac{e^{sim(\boldsymbol{z}^{spe}_{i},\boldsymbol{z}^{tem}_{i})/\tau}}{\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}e^{sim(\boldsymbol{z}^{spe}_{i},\boldsymbol{z}^{spe}_{j})/\tau}+\sum_{j=1}^{N}e^{sim(\boldsymbol{z}^{spe}_{i},\boldsymbol{z}^{tem}_{j})/\tau}},$

### v2

@@ 3.3 Data Acquisition and Preprocessing
Figure 4: The structure of our proposed TSformer-SA. The $F(\cdot)$ represents the token score function, while $\mathcal{L}_{cls}$ and $\mathcal{L}_{multi\mbox{-}view}$ respectively denote the cross-entropy loss and the multi-view consistency loss. The inputs consist of EEG temporal signals representing the temporal view and the spectrogram images representing the spectral view. The feature extractor tokenizes the inputs and extracts the view-specific features. Subsequently, the cross-view interaction module extracts the common features from both views and the fusion module fuses the two-view features for classification. The above three modules are trained during the pre-training stage and only the subject-specific adapter is trained in the fine-tuning stage.
@@ 4 Method
In this section, we introduce the Temporal-Spectral fusion transformer with Subject-specific Adapter (TSformer-SA) for RSVP decoding. As shown in Fig. 4, TSformer-SA is a symmetrical dual-stream Transformer comprising a feature extractor, a cross-view interaction module, a fusion module, and a subject-specific adapter. The inputs of the model are EEG temporal signals as the temporal view and the spectrogram image of each channel computed using CWT as the spectral view. First, the feature extractor employs a slice embedding layer to tokenize the two-view input and an encoder layer to extract view-specific features. Subsequently, the cross-view interaction module employs cross-attention and token fusion mechanisms to realize information transfer between the two views. Simultaneously, we introduce a multi-view consistency loss for the common feature extraction between two views by maximizing the similarity of different view features from the same sample. Then, an attention-based fusion module achieves the feature fusion between two views to obtain comprehensive features for classification.
The optimization strategy of TSformer-SA consists of the pre-training stage and the adapter-based fine-tuning stage. In the pre-training stage, multi-view consistency loss and cross-entropy loss are used to optimize TSformer on data from existing subjects. Then the proposed subject-specific adapter is fine-tuned on new subjects’ data to rapidly transfer the learned general patterns of the pre-trained model to the data distribution of new subjects. To facilitate reading, the frequently used symbols and their definitions are listed Table 1.
Table 1: The frequently used symbols and their definitions.
Symbol Shape Definition
$\boldsymbol{S}_{tem}$ $\mathbb{R}^{1\times C\times T}$ EEG temporal signals, with $C$ EEG channels and $T$ time points.
$\boldsymbol{S}_{spe}$ $\mathbb{R}^{S\times C\times T}$ EEG spectrogram images, with $S$ scale.
$\boldsymbol{X}_{tem},\boldsymbol{X}_{spe}$ $\mathbb{R}^{n\times d}$ Temporal and spectral token sequences, with $n$ tokens.
$\boldsymbol{x}_{i}^{tem}$, $\boldsymbol{x}_{i}^{spe}$ $\mathbb{R}^{d}$ The $i$-th temporal token and spectral token in $\boldsymbol{X}_{tem}$ and $\boldsymbol{X}_{spe}$ respectively, with $d$ dimension.
$\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i}$ $\mathbb{R}^{3d}$ The feature vectors of the temporal view and spectral view from the $i$-th sample used in the multi-view consistency loss.
$\boldsymbol{W}_{Q},\boldsymbol{W}_{K},\boldsymbol{W}_{V}$ $\mathbb{R}^{d\times d}$ Learnable parameter matrix in attention mechanisms.
$\boldsymbol{W}_{score}$ $\mathbb{R}^{d\times d}$ Learnable parameter matrix in token score function.
$F(\cdot)$ Token score function.
[…]
@@ 4.1 Feature Extractor
The feature extractor includes a slice embedding layer and an encoder layer. Since the difference between the P300 signals induced by target images and the harmonic signals induced by non-target images in the RSVP task is concentrated along the time dimension, the preprocessed EEG temporal signals and spectrogram images are initially divided into non-overlapping slices along the time dimension. To notice more valuable time slice information, we adopt the self-attention mechanism within each view to perceive the global time dependencies of EEG temporal signals and spectrogram images.
@@
The inputs of the slice embedding layer are EEG temporal signals ($\boldsymbol{S}_{tem}\in\mathbb{R}^{1\times C\times T}$) and spectrogram images of each channel ($\boldsymbol{S}_{spe}\in\mathbb{R}^{S\times C\times T}$), where $C$ denotes the number of channels, $T$ denotes the number of sampling points, and $S$ denotes the size of scale in CWT. We slice and reshape the input signals $\boldsymbol{S}_{tem}$ and $\boldsymbol{S}_{spe}$ into the sequence of flattened slices $\left[\boldsymbol{s}_{1}^{tem},\boldsymbol{s}_{2}^{tem},\cdots,\boldsymbol{s}_{n}^{tem}\right]$ and $\left[\boldsymbol{s}_{1}^{spe},\boldsymbol{s}_{2}^{spe},\cdots,\boldsymbol{s}_{n}^{spe}\right]$ along the time dimension, respectively. Each $\boldsymbol{s}_{i}^{tem}\in\mathbb{R}^{(C\times t)}$ and $\boldsymbol{s}_{i}^{spe}\in\mathbb{R}^{(S\times C\times t)}$ respectively represents brain temporal and spectral activities over a period of time, where $t$ is the slice length and $n=[T/t]$ is the number of slices. Then, the slices are projected to $d$ dimensions, where the output of these temporal and spectral slice embeddings are referred to as temporal tokens $\boldsymbol{X}_{tem}\in\mathbb{R}^{n\times d}$ and spectral tokens $\boldsymbol{X}_{spe}\in\mathbb{R}^{n\times d}$, respectively. The process can be formulated as follows:
$\boldsymbol{X}_{tem}=\left[\boldsymbol{s}_{1}^{tem},\boldsymbol{s}_{2}^{tem},\cdots,\boldsymbol{s}_{n}^{tem}\right]^{T}\boldsymbol{W}_{pos}^{(1)}+\boldsymbol{E}_{pos}^{(1)},$ (1)
$\boldsymbol{X}_{spe}=\left[\boldsymbol{s}_{1}^{spe},\boldsymbol{s}_{2}^{spe},\cdots,\boldsymbol{s}_{n}^{spe}\right]^{T}\boldsymbol{W}_{pos}^{(2)}+\boldsymbol{E}_{pos}^{(2)},$ (2)
where $\boldsymbol{W}_{pos}^{(i)}\in\mathbb{R}^{(C\times t)\times d}$ ($i=1,2$) is the parameter matrix and $\boldsymbol{E}_{pos}^{(i)}\in\mathbb{R}^{n\times d}$ ($i=1,2$) is the learnable positional parameter matrix added to the slice embeddings to retain positional information. The resulting sequence of embedding vectors serves as the input to the encoder layers.
@@
The encoder layer comprises two sub-layers: a Multi-Head Self-Attention (MHSA) layer capturing global temporal dependencies within input tokens, and a position-wise Feed-Forward Network (FFN) layer with a hidden layer consisting of $h\times d$ dimensions extracting feature representations, where $h$ denotes the number of heads. All sub-layers utilize a residual connection and layer normalization to enhance the scalability of the model. To reduce model parameters and enable the subsequent view interaction more reasonable and feasible, we share the encoder parameters between the two streams to project the features into the same feature space. The outputs of the feature extractor are temporal tokens ($\boldsymbol{X}_{tem}$) and spectral tokens ($\boldsymbol{X}_{spe}$).
@@ 4.2.1 Cross-View Interaction Module
The cross-view interaction module consists of a cross-attention layer and a token fusion layer. To capitalize on the complementary information between two views and extract common features, the cross-attention mechanism is first employed to focus on EEG information transfer between temporal and spectral views and adaptively learn inter-view common representations.
@@
The input of the cross-attention layer are token sequences of two views: $\boldsymbol{X}_{tem}=[\boldsymbol{x}^{tem}_{1},\boldsymbol{x}^{tem}_{2},\cdots,\boldsymbol{x}^{tem}_{n}]^{T}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{X}_{spe}=[\boldsymbol{x}^{spe}_{1},\boldsymbol{x}^{spe}_{2},\cdots,\boldsymbol{x}^{spe}_{n}]^{T}\in\mathbb{R}^{n\times d}$, where $\boldsymbol{x}^{tem}_{i}\in\mathbb{R}^{d}$ and $\boldsymbol{x}^{spe}_{i}\in\mathbb{R}^{d}$ denote the $i$-th token in temporal and spectral token sequences, respectively. The linear projection layer maps input token sequences to sequential vectors and then both views mutually extract common features from each other based on the similarity between their tokens. We utilize a multi-head attention mechanism. To concisely represent the model inference process, we illustrate the single-head attention mechanism as an example. The process can be formulated as follows:
$\boldsymbol{X}_{tem}^{CA}=\boldsymbol{X}_{tem}+\sigma(\frac{\boldsymbol{X}_{tem}\boldsymbol{W}_{Q}^{(1)}{\boldsymbol{W}_{K}^{(1)}}^{T}\boldsymbol{X}_{spe}^{T}}{\sqrt{d_{k}}})\boldsymbol{X}_{spe}\boldsymbol{W}_{V}^{(1)},$ (3)
$\boldsymbol{X}_{spe}^{CA}=\boldsymbol{X}_{spe}+\sigma(\frac{\boldsymbol{X}_{spe}\boldsymbol{W}_{Q}^{(2)}{\boldsymbol{W}_{K}^{(2)}}^{T}\boldsymbol{X}_{tem}^{T}}{\sqrt{d_{k}}})\boldsymbol{X}_{tem}\boldsymbol{W}_{V}^{(2)},$ (4)
where $\boldsymbol{X}_{tem}^{CA}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{X}_{spe}^{CA}\in\mathbb{R}^{n\times d}$ are the temporal and spectral view output of the cross-attention layer respectively. The linear matrices $\boldsymbol{W}_{Q}^{(i)}\in\mathbb{R}^{d\times d}$, $\boldsymbol{W}_{K}^{(i)}\in\mathbb{R}^{d\times d}$, and $\boldsymbol{W}_{V}^{(i)}\in\mathbb{R}^{d\times d}$ ($i=1,2$) are learnable parameter matrices. The $\sigma(\cdot)$ represents the softmax function.
As the low signal-to-noise ratio of the single sample EEG signal may lead to redundant tokens embedded from some EEG time periods, it is crucial to dynamically identify and replace these uninformative tokens with features pertinent to decoding tasks. We propose a token score function to measure the importance of each token within each view and then apply token fusion [60] to selectively prune uninformative tokens and replace each pruned token using projected features at the same time period from alternative views.
For the updating of temporal tokens $\boldsymbol{X}_{tem}^{CA}$, we assess the importance of token $\boldsymbol{x}^{tem}_{i}$ using a scoring function denoted as $f(\cdot)$:
@@
$f(\boldsymbol{x}^{tem}_{i})=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\sigma\left(\dfrac{\boldsymbol{X}_{tem}^{CA}\boldsymbol{W}_{score}{\boldsymbol{X}_{tem}^{CA}}^{T}}{\sqrt{d}}\right)_{(j,i)},$ (5)
@@
where $\boldsymbol{W}_{score}\in\mathbb{R}^{d\times d}$ is a learnable parameter matrix. The scoring function represents the accumulation of attention weights from other tokens towards $\boldsymbol{x}^{tem}_{i}$ in the self-attention matrix of temporal tokens $\boldsymbol{X}_{tem}^{CA}$. Tokens that receive higher attention can be considered to contain more information. On the basis of $f(\cdot)$, we can establish the scoring function of temporal tokens as $F(\cdot)$, which can be formulated as:
$F(\boldsymbol{X}_{tem}^{CA})=\left[f(\boldsymbol{x}^{tem}_{1})\cdot\boldsymbol{1},f(\boldsymbol{x}^{tem}_{2})\cdot\boldsymbol{1},\cdots,f(\boldsymbol{x}^{tem}_{n})\cdot\boldsymbol{1}\right],$ (6)
@@
$\boldsymbol{X}_{tem}^{TF}=\dfrac{1}{2}\left[\boldsymbol{X}_{tem}^{CA}+\boldsymbol{X}_{spe}^{CA}\odot\mathbb{I}_{(F(\boldsymbol{X}_{tem}^{CA})<\theta_{tem})}\odot\mathbb{I}_{(F(\boldsymbol{X}_{spe}^{CA})>\theta_{spe})}\right],$ (7)
$\boldsymbol{X}_{spe}^{TF}=\dfrac{1}{2}\left[\boldsymbol{X}_{spe}^{CA}+\boldsymbol{X}_{tem}^{CA}\odot\mathbb{I}_{(F(\boldsymbol{X}_{spe}^{CA})<\theta_{spe})}\odot\mathbb{I}_{(F(\boldsymbol{X}_{tem}^{CA})>\theta_{tem})}\right],$ (8)
where the threshold $\theta_{tem}$ and $\theta_{spe}$ are equal to the median of $F(\boldsymbol{X}_{tem}^{CA})$ and $F(\boldsymbol{X}_{spe}^{CA})$, respectively. The $\boldsymbol{X}_{tem}^{TF},\boldsymbol{X}_{spe}^{TF}\in\mathbb{R}^{n\times d}$ denote the temporal and spectral view output of the token fusion and $\mathbb{I}_{(\cdot)}$ is an indicator asserting the subscript condition, which outputs a mask tensor with element 0 or 1. The operator $\odot$ denotes the element-wise multiplication. Finally, we use a self-attention layer to extract the global features of tokens with added supplementary information.
@@ 4.2.2 Multi-View Consistency Loss
The temporal and spectral features respectively represent the temporal and spectral activities of EEG signals and there are great differences between the two view features. To extract common features from both views to enhance decoding performance, we design a multi-view consistency loss that narrows the distance between the temporal and spectral features from the same sample based on contrastive learning [61].
Firstly, a convolution layer with four convolution kernels, each of size $(16,d/8)$ and stride $(16,d/8)$, is utilized to map the token sequences of both views ($\boldsymbol{X}_{tem}^{TF}$, $\boldsymbol{X}_{spe}^{TF}$) to features $\{\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i}\}_{i=1}^{N}$ in a shared feature space, where $\boldsymbol{z}^{tem}_{i}$ and $\boldsymbol{z}^{spe}_{i}$ are the feature vectors of the temporal view and spectral view from the $i$-th sample, respectively. The $N$ denotes the batch size. Feature similarity is assessed using the cosine distance $sim(\boldsymbol{u},\boldsymbol{v})=\boldsymbol{u}^{T}\boldsymbol{v}/\|\boldsymbol{u}\|_{2}\|\boldsymbol{v}\|_{2}$, and the loss function with respect to positive pairs $(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})$ is defined as:
$g(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})=\frac{e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})/\tau}}{\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{tem}_{j})/\tau}+\sum_{j=1}^{N}e^{sim(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{j})/\tau}},$ (9)
$\mathcal{L}_{i}=\log\left[g(\boldsymbol{z}^{tem}_{i},\boldsymbol{z}^{spe}_{i})\right]+\log\left[g(\boldsymbol{z}^{spe}_{i},\boldsymbol{z}^{tem}_{i})\right],$ (10)